What pressure is exerted by a mixture of 2.10 g of H2 and 8.700 g of N2 at 273°C in a 10.0 L container?
Answer #1

Answer:-

Given:-

volume of container (V) = 10.0 L

temperature (T) = 273 0C = 273 + 273 = 546 K

wt. of H2 gas = 2.10 g

wt. of N2 gas = 8.700 g

partial pressure of H2 gas (PH2) = ?

partial pressure of N2 gas (PN2) = ?

total pressure exerted by a mixture of H2 and N2 gases in the container (Ptotal) =?

As we know that

molar mass of H2 gas = 2 \times molar mass of H

molar mass of H2 gas = 2 \times 1

molar mass of H2 gas = 2 g /mol

similarly

molar mass of N2 gas = 2 \times molar mass of N

molar mass of N2 gas = 2 \times 14

molar mass of N2 gas = 28 g /mol

So

no. of moles of H2 gas (nH2) = wt. of H2 gas / molar mass of H2 gas

no. of moles of H2 gas (nH2) = 2.10 g / 2 g /mol

no. of moles of H2 gas (nH2) = 1.05 mol

similarly

no. of moles of N2 gas (nN2) = wt. of N2 gas / molar mass of N2 gas

no. of moles of N2 gas (nN2) =  8.700 g / 28 g /mol

no. of moles of N2 gas (nN2) = 0.311 mol

As we know that

gas constant (R) = 0.08206 L.atm.K-1mol-1

So according to formula

PV = nRT

P = nRT / V

therefore

partial pressure of H2 gas (PH2) = no. of moles of H2 gas (nH2) \times gas constant (R) \times temperature (T) / volume of container (V)

partial pressure of H2 gas (PH2) = 1.05 mol \times 0.08206 L.atm.K-1mol-1\times 546 K / 10.0 L

partial pressure of H2 gas (PH2) = 47.044998‬ L.atm / 10.0 L

partial pressure of H2 gas (PH2) = 4.70 atm

similarly

partial pressure of N2 gas (PN2) = no. of moles of N2 gas (nH2) \times gas constant (R) \times temperature (T) / volume of container (V)

partial pressure of N2 gas (PN2) = 0.311 mol \times 0.08206 L.atm.K-1mol-1\times 546 K / 10.0 L

partial pressure of N2 gas (PN2) = 13.93428036‬ L.atm / 10.0 L

partial pressure of N2 gas (PN2) = 1.39 atm

Also we know that according to the Dalton's Law of partial pressure " the total pressure of mixture of gases is equal to the sum of all the partial pressures of gases present in the mixture".

therefore

total pressure exerted by a mixture of H2 and N2 gases in the container (Ptotal) = partial pressure of H2 gas (PH2) + partial pressure of N2 gas (PN2)

total pressure exerted by a mixture of H2 and N2 gases in the container (Ptotal) = 4.70 atm  + 1.39 atm

total pressure exerted by a mixture of H2 and N2 gases in the container (Ptotal) = 6.09 atm (i.e the answer)